Peter is one-fifth of his fathers age. Their combined age is 48. Write a linear equation to represent this information. Solve the linear equation. How old is Peter's Father and how old is Peter?
Answer by addingup(3677)

Peter: P
Father: F
You can substitute with x and y if required.
:
P = 1/5F
P+F = 48 substitute for P
1/5F+F = 48
Multiply both sides times 5:
F+5F = 240
6F = 240
F = 240/6 = 40 The father is 40
And Peter: 40*1/5 = 40/5 = 8 Peter is 8